OpenStudy (anonymous):

consider the function represented by 9x+3y=12 with x as the independent variable. how can this function be written using function notation?

OpenStudy (ehsan18):

take out y from the equation and put it equal to f(x) for example 3y=12-9x y=(12-9x)/3 f(x)=(12-9x)/3

OpenStudy (anonymous):

On the bottom of the question is says f(x)=________x+4 so I don't really understand still what would go in the blank